# DialUp! Modeling the Language Continuum by Adapting Models to Dialects and Dialects to Models

### v1

@@ Abstract
Most of the world’s languages and dialects are low-resource, and lack support in mainstream machine translation (MT) models. However, many of them have a closely-related high-resource language (HRL) neighbor, and differ in linguistically regular ways from it. This underscores the importance of model robustness to dialectical variation and cross-lingual generalization to the HRL dialect continuum. We present DialUp, consisting of a training-time technique for adapting a pretrained model to dialectical data (MD), and an inference-time intervention adapting dialectical data to the model expertise (DM). MD induces model robustness to potentially unseen and unknown dialects by exposure to synthetic data exemplifying linguistic mechanisms of dialectical variation, whereas DM treats dialectical divergence for known target dialects. These methods show considerable performance gains for several dialects from four language families, and modest gains for two other language families. We also conduct feature and error analyses, which show that language varieties with low baseline MT performance are more likely to benefit from these approaches.111https://github.com/niyatibafna/dialectical-robustness-mt/
@@ 1 Introduction
The roughly 7000 languages in the world can be grouped into a few hundred families. Closely-related languages (CRLs) and dialects within a family largely exhibit continuous and structured differences along phonological, morphological, and lexical dimensions, rather than being discrete monoliths. Hovy and Purschke (2018); Bergman and Diab (2022). Not only is it unfeasible to collect training data for every lect, but language is fluid and dynamic across and between dialect boundaries, calling for principled general approaches to dialectical variation.
@@
We first aim to adapt the model to the language continuum, or model-to-data (MD) adaptation (yellow arrows in Figure 1). The goal is robustness to unseen CRLs (both known CRLs absent in pre-training data and yet undocumented CRL varieties), using only HRL bitext. We do this by simulating dialectical variation patterns over HRL fine-tuning text, teaching the model to generalise to realistic variants such as actual CRL inputs.
@@
Note that cross-lingual generalization can be viewed as a train-test data mismatch problem: we want our models to work on dialectical data which differs in various ways from the training data of the model (HRL data). Our two approaches above then represent two broad paradigms for this general problem in machine learning (see Figure 1). Model-to-data adaptation has roots in existing cross-lingual transfer approaches, which approximate CRLs using an HRL, relying on linguistic similarities for transfer; however, this does not train the model to handle dialectical divergences. MD innovates on this by synthetically approximating such divergences. By approximating language continua, we may avoid pressing questions raised by discrete language paradigms, e.g. which transfer language to use and how much training data to collect (Dalmia et al., 2018). DM adaptation has roots in prior approaches to approximate a model’s expected domain in data (Nie et al., 2023).
@@
Gains from DM adaptation over baselines for certain families and CRLs (yielding up to mean BLEU for Indic dialects), showing for the first time that adapting dialectical function words is more beneficial than adapting content words
Evidence that MD and DM combine advantageously, and provide a recipe for increasing the flexibility of existing MT models to general dialectical variation.
@@ 2.1 Model-to-data (MD)
In this paradigm, given an HRL-proficient model and bitext in that HRL, we adapt the model to unseen CRLs. We generate varied artificial CRLs of the HRL by simulating mechanisms of dialectical variation over the HRL bitext, and fine-tune on the resulting synthetic data to induce model robustness to dialectical variation.
@@ 4 Results
MD gives mean gains over best baselines for all language families, demonstrating the efficacy of linguistically-motivated (as opposed to random) synthetic data augmentation. Table 4 shows examples of improvements on cognate words, including new inflections and function words. We often see that while off-the-shelf models transliterated them or left them as are instead of translating, MD-tuned models decode them correctly.101010See full-length examples in Appendix G. DM introduces gains for many low-performing varieties—up to BLEU for Aya-23 on Sundanese and BLEU for M2M on Chhattisgarhi (hne)—where replacing dialectal words with HRL equivalents is helpful. However, it is also damaging for several, especially high-performing CRLs like French (fra), Spanish (spa), and Tagalog (tgl). Conceivably, if the base model is already proficient in a CRL, DM code-switching introduces counterproductive unnaturalness. Since DM is an inference-time method, practitioners may activate or deactivate it according to language needs. See Table 5 for examples of DM-func input and output, demonstrating the brittleness of baseline models on dialectical function words, as well as the impact of treating them.
@@ 5 Discussion
Measurements of HRL-CRL relatedness had the second-highest forest feature importance for both models. Though our measurements of HRL-CRL relatedness do not correlate with BLEU score improvement, we suspect there may be a non-linear relationship between dev-set character-F score and method effectiveness. When plotted in Figure 6, the points suggest the contour of a downward-facing parabola, indicating that MD is most effective for CRLs that are not too close but not too far from the HRL, as we hypothesized. This has an intuitive interpretation: CRLs that are too close may either already be well-performing, or may not display enough dialectical divergence to benefit from DialUp, whereas CRLs that are too far may have a higher amount of non-cognate divergences from the HRL that DialUp does not help with. However, note that language family is possibly confounded with character-F score, with only Austronesian CRLs covering a wide range on our observed contour; it would be difficult to disentangle the effect of language family without many more languages.
@@ 7 Conclusion
We provide results on one representative model of the traditional NMT as well as LLM paradigm, choosing an appropriate model with respect to its training languages given our goal of handling unseen and unknown target dialects, and available evaluation data. Naturally, our method should be tested on a variety of models from each paradigm to gauge its usefulness as a general data augmentation technique for achieving dialectical robustness.

### v2

@@ Abstract
Most of the world’s languages and dialects are low-resource, and lack support in mainstream machine translation (MT) models. However, many of them have a closely-related high-resource language (HRL) neighbor, and differ in linguistically regular ways from it. This underscores the importance of model robustness to dialectal variation and cross-lingual generalization to the HRL dialect continuum. We present DialUp, consisting of a training-time technique for adapting a pretrained model to dialectal data (MD), and an inference-time intervention adapting dialectal data to the model expertise (DM). MD induces model robustness to potentially unseen and unknown dialects by exposure to synthetic data exemplifying linguistic mechanisms of dialectal variation, whereas DM treats dialectal divergence for known target dialects. These methods show considerable performance gains for several dialects from four language families, and modest gains for two other language families. We also conduct feature and error analyses, which show that language varieties with low baseline MT performance are more likely to benefit from these approaches.111https://github.com/niyatibafna/dialup/
@@ 1 Introduction
The roughly 7000 languages in the world can be grouped into a few hundred families. Closely-related languages (CRLs) and dialects within a family largely exhibit continuous and structured differences along phonological, morphological, and lexical dimensions, rather than being discrete monoliths. Hovy and Purschke (2018); Bergman and Diab (2022). Not only is it unfeasible to collect training data for every lect, but language is fluid and dynamic across and between dialect boundaries, calling for principled general approaches to dialectal variation.
@@
We first aim to adapt the model to the language continuum, or model-to-data (MD) adaptation (yellow arrows in Figure 1). The goal is robustness to unseen CRLs (both known CRLs absent in pre-training data and yet undocumented CRL varieties), using only HRL bitext. We do this by simulating dialectal variation patterns over HRL fine-tuning text, teaching the model to generalise to realistic variants such as actual CRL inputs.
@@
Note that cross-lingual generalization can be viewed as a train-test data mismatch problem: we want our models to work on dialectal data which differs in various ways from the training data of the model (HRL data). Our two approaches above then represent two broad paradigms for this general problem in machine learning (see Figure 1). Model-to-data adaptation has roots in existing cross-lingual transfer approaches, which approximate CRLs using an HRL, relying on linguistic similarities for transfer; however, this does not train the model to handle dialectal divergences. MD innovates on this by synthetically approximating such divergences. By approximating language continua, we may avoid pressing questions raised by discrete language paradigms, e.g. which transfer language to use and how much training data to collect (Dalmia et al., 2018). DM adaptation has roots in prior approaches to approximate a model’s expected domain in data (Nie et al., 2023).
@@
Gains from DM adaptation over baselines for certain families and CRLs (yielding up to mean BLEU for Indic dialects), showing for the first time that adapting dialectal function words is more beneficial than adapting content words
Evidence that MD and DM combine advantageously, and provide a recipe for increasing the flexibility of existing MT models to general dialectal variation.
@@ 2.1 Model-to-data (MD)
In this paradigm, given an HRL-proficient model and bitext in that HRL, we adapt the model to unseen CRLs. We generate varied artificial CRLs of the HRL by simulating mechanisms of dialectal variation over the HRL bitext, and fine-tune on the resulting synthetic data to induce model robustness to dialectal variation.
@@ 4 Results
MD gives mean gains over best baselines for all language families, demonstrating the efficacy of linguistically-motivated (as opposed to random) synthetic data augmentation. Table 4 shows examples of improvements on cognate words, including new inflections and function words. We often see that while off-the-shelf models transliterated them or left them as are instead of translating, MD-tuned models decode them correctly.101010See full-length examples in Appendix G. DM introduces gains for many low-performing varieties—up to BLEU for Aya-23 on Sundanese and BLEU for M2M on Chhattisgarhi (hne)—where replacing dialectal words with HRL equivalents is helpful. However, it is also damaging for several, especially high-performing CRLs like French (fra), Spanish (spa), and Tagalog (tgl). Conceivably, if the base model is already proficient in a CRL, DM code-switching introduces counterproductive unnaturalness. Since DM is an inference-time method, practitioners may activate or deactivate it according to language needs. See Table 5 for examples of DM-func input and output, demonstrating the brittleness of baseline models on dialectal function words, as well as the impact of treating them.
@@ 5 Discussion
Measurements of HRL-CRL relatedness had the second-highest forest feature importance for both models. Though our measurements of HRL-CRL relatedness do not correlate with BLEU score improvement, we suspect there may be a non-linear relationship between dev-set character-F score and method effectiveness. When plotted in Figure 6, the points suggest the contour of a downward-facing parabola, indicating that MD is most effective for CRLs that are not too close but not too far from the HRL, as we hypothesized. This has an intuitive interpretation: CRLs that are too close may either already be well-performing, or may not display enough dialectal divergence to benefit from DialUp, whereas CRLs that are too far may have a higher amount of non-cognate divergences from the HRL that DialUp does not help with. However, note that language family is possibly confounded with character-F score, with only Austronesian CRLs covering a wide range on our observed contour; it would be difficult to disentangle the effect of language family without many more languages.
@@ 7 Conclusion
We provide results on one representative model of the traditional NMT as well as LLM paradigm, choosing an appropriate model with respect to its training languages given our goal of handling unseen and unknown target dialects, and available evaluation data. Naturally, our method should be tested on a variety of models from each paradigm to gauge its usefulness as a general data augmentation technique for achieving dialectal robustness.